mathslover (mathslover):

Hi i am confused in a problem : a^2-b^2=(a+b)(a-b) right? now i have to prove that a^2-b^2=(a+b)(a-b) I know that RHS can be expanded and this formula can be proved but m confused how to do that by just taking LHS ?? Help please

mathslover (mathslover):

I hope that you all got what I meant to say..

OpenStudy (anonymous):

\[a^2-b^2+ab-ab=a^2+ab-b^2-ab=a(a+b)-b(a+b)\]does that help?

OpenStudy (anonymous):

add ab and subtract ab.......... and factor it

mathslover (mathslover):

oh i got it thanks a lot @mukushla and @sauravshakya thanks again.. that was just a small confusion that is resolved now.. thanks

OpenStudy (anonymous):

Welcome...

OpenStudy (anonymous):

yw :)

hero (hero):

What the...@mathslover, I assumed you already knew this kinda stuff.

OpenStudy (saifoo.khan):

Nice time-passing methods.

mathslover (mathslover):

@Hero i knew that stuff but not about that LHS .. sometimes "genius" persons also make mistake.. but I am not so genius. It was not time pass @saifoo.khan

OpenStudy (amistre64):

since we can expand the RHS, why would we need to reprove it using the LHS?

mathslover (mathslover):

It was my curiousness ...

mathslover (mathslover):

for example we have proof for (a+b)^2=a^2+b^2+2ab we can proove that by using LHS also

mathslover (mathslover):

\[{(a+b)^2=(a+b)(a+b)=a(a+b)+b(a+b)\\=a^2+ab+ba+b^2=a^2+2ab+b^2}\]

OpenStudy (amistre64):

(a+b)(a+b)

mathslover (mathslover):

right

OpenStudy (amistre64):

im not sure if i would call these proofs, moreso the results of using appropriate mathings

mathslover (mathslover):

good point "sir"

OpenStudy (amistre64):

if you can work out one side to get to the other; the proof of the other side is just the reversing of the steps

mathslover (mathslover):

Yep

OpenStudy (shubhamsrg):

a^2 - b^2 = a^2 + b^2 - 2b^2 + 2ab - 2ab = (a+b)^2 - 2b(a+b) =>(a+b)(a-b) there must be many more methods !